# Nonisotropic Gaussian Diffusion for Realistic 3D Human Motion Prediction

### v2

@@ 4.1 Nonisotropic Gaussian Diffusion
where ${\gamma}_{t}=1-t/T$ defines a monotonically decreasing scheduler. Detailed derivation and alternative scheduler formulation explored in early stages are in Appendix A.
@@
$L_{\mathrm{diff}}({\boldsymbol{x}}_{\boldsymbol{\theta}},{\boldsymbol{x}}_{0},t):=\bar{\alpha}_{t}\|\bar{\boldsymbol{\Lambda}}_{t}^{-1/2}\boldsymbol{U}^{\top}({\boldsymbol{x}}_{\boldsymbol{\theta}}-{\boldsymbol{x}}_{0})\|^{2}.$ (8)
@@ 4.3 SkeletonDiffusion
In latent space, the denoising network $\mathrm{{g}}$ learns via our nonisotropic diffusion formulation to denoise true latent variables $\boldsymbol{z}$ conditioned on past observations $\mathrm{{e}}(\textbf{X})$. Instead of predicting the noise ${\boldsymbol{\epsilon}}_{t}$ [31, 64], we directly approximate the true latent code ${\boldsymbol{x}}_{0}:=\boldsymbol{z}$ [63, 5] as ${\boldsymbol{x}}_{\boldsymbol{\theta}}=\mathrm{{g}}({\boldsymbol{x}}_{t},\mathrm{{e}}(\textbf{X}),t)$. To implicitly enforce diversity [5, 69, 29], we relax Eq. (8) by sampling $k=50$ predictions at each iteration and backpropagating the gradient only through the sample closest to the GT:
@@
Instead of choosing the sample that minimizes the diffusion loss [5, 69, 29], we choose the prediction that minimizes the reconstruction loss Eq. 11, finding that this benefits diversity in the generated ensemble. At inference, we denoise multiple latent codes ${\boldsymbol{x}}_{\theta}$ according to our reverse formulation LABEL:eq:reverse_diff. The generated latent codes are then decoded into future predictions $\tilde{\textbf{Y}}$.
@@ E.3 On the Convergence of Nonisotropic Diffusion
E.4 Ablations of SkeletonDiffusion
In Tab. 7, we report the ablations discussed in Sec. 5.2. We compare the effect of TG-Attention layers on isotropic diffusion ($\boldsymbol{\Sigma}_{N}=I$ and ${\gamma}_{t}=0$) and analyze nonisotropic diffusion with our covariance reflecting joint connections $\boldsymbol{\Sigma}_{N}$ (Eq. 4) in the variant where ${\gamma}_{t}=1$ (as in Eq. 3) and our blending with the scheduler ${\gamma}_{t}$ (Eq. 5). Further experiments, such as fine-tuning the encoder responsible for embedding the past observation or representing motion data via the Discrete Cosine Transform (DCT) [16] are reported in Tab. 8. From the low precision results of the latter experiment and referring to Tab. 1, we speculate that while DCT seems suitable for transformer-based diffusion models operating in input space [69, 16], extracting features directly from Euclidean motion space seems a better choice for latent diffusion models (BeLFusion [5] and our method).
Appendix F Additional Experiments
@@
Table 12: Model footprint for a single H36M inference (RTX 6000)
F.1 Diversity and Body Realism

### v3

@@ 4.1 Nonisotropic Gaussian Diffusion
where ${\gamma}_{t}$ defines a cosine scheduler with opposite behavior to ${\alpha}_{t}$. Detailed derivation and alternative scheduler formulation explored in early stages are in Appendix A.
@@
$L_{\mathrm{diff}}({\boldsymbol{x}}_{\boldsymbol{\theta}},{\boldsymbol{x}}_{0},t):=\bar{\alpha}_{t}\|\bar{\boldsymbol{\Lambda}}_{t}^{-1/2}\boldsymbol{U}^{\top}({\boldsymbol{x}}_{\boldsymbol{\theta}}-{\boldsymbol{x}}_{0})\|.$ (8)
@@ 4.3 SkeletonDiffusion
In latent space, the denoising network $\mathrm{{g}}$ learns via our nonisotropic diffusion formulation to denoise true latent variables $\boldsymbol{z}=\mathrm{{e}}(\textbf{Y})$ conditioned on past observations $\mathrm{{e}}(\textbf{X})$. Instead of predicting the noise ${\boldsymbol{\epsilon}}_{t}$ [31, 64], we directly approximate the true latent code ${\boldsymbol{x}}_{0}:=\boldsymbol{z}$ [63, 5] as ${\boldsymbol{x}}_{\boldsymbol{\theta}}=\mathrm{{g}}({\boldsymbol{x}}_{t},\mathrm{{e}}(\textbf{X}),t)$. To implicitly enforce diversity [5, 69, 29], we relax Eq. (8) by sampling $k=50$ predictions at each iteration and backpropagating the gradient only through the sample closest to the GT:
@@
Instead of choosing the sample that minimizes the diffusion loss [5, 69, 29], we choose the prediction that minimizes the reconstruction loss Eq. 11, finding that this benefits diversity in the generated ensemble (Sec. E.4). At inference, we denoise multiple latent codes ${\boldsymbol{x}}_{\theta}$ according to our reverse formulation LABEL:eq:reverse_diff. The generated latent codes are then decoded into future predictions $\tilde{\textbf{Y}}$.
@@ E.3 On the Convergence of Nonisotropic Diffusion
E.4 Ablations of SkeletonDiffusion
In Tab. 7, we report the ablations discussed in Sec. 5.2. We compare the effect of TG-Attention layers on isotropic diffusion ($\boldsymbol{\Sigma}_{N}=I$ and ${\gamma}_{t}=0$) and analyze nonisotropic diffusion with our covariance reflecting joint connections $\boldsymbol{\Sigma}_{N}$ (Eq. 4) in the variant where ${\gamma}_{t}=1$ (as in Eq. 3) and our blending with the scheduler ${\gamma}_{t}$ (Eq. 5).
In Tab. 8, we report further experiments, such as fine-tuning the encoder responsible for embedding the past observation (Ours+Past) or representing motion data via the Discrete Cosine Transform (DCT) [16]. From the low precision results of the DCT experiment (Ours+DCT) and referring to Tab. 1, we speculate that while DCT seems suitable for transformer-based diffusion models operating in input space [69, 16], extracting features directly from Euclidean motion space seems a better choice for latent diffusion models (BeLFusion [5] and our method).
Diffusion Training Objective and k-Relaxation
In the same table Tab. 8, we also ablate the relaxation of the diffusion objective described in Sec. 4.3 Eq. 12 with three experiments: (1) by not relaxing the diffusion loss during training i.e. setting $k=1$ and backpropagating the loss through the first and only sample; (2) by sampling $k=50$ times and backpropagating through the generated sample that is most similar to the ground truth in latent space; and (3) doing so with $k=50$ but in motion space - which is the choice for SkeletonDiffusion. By looking at the result for the first two cases, we see that $k=1$ generates considerably less diverse futures, confirming the detailed investigation of BeLFusion [5] on how increasing $k$ leads to higher diversity. In our case, setting $k=50$ does not only lead to a double as high diversity score, but also improves precision and realism. We thus believe relaxation to be a strong guidance towards the target distribution, particularly for latent space models. Intuitively, it allows for better coverage of the different future modes of the real data distribution. In SkeletonDiffusion, we go a step further and choose the sample to backpropagate the loss not via similarity in latent space, but in motion space. This design choice improves diversity by more than 10% (!). Intuitively, in the early stage of the diffusion training, the denoiser generates coarse latent codes, whose similarity to the ground truth embedding may ’erroneously’ not reflect the effective perceptual similarity in motion or input space, leading to suboptimal training. We note, though, that this choice leads to considerably increased training time. Due to the recurrent nature of the decoder, decoding the $k=50$ generated samples in motion space leads to more than doubled training time: the diffusion network of SkeletonDiffusion trains in 5 days, while the version with $k=1$ in a single day, and $k=50$ in latent space in .ca a day and half.
Appendix F Additional Experiments
F.1 Diversity and Body Realism
@@
Table 12: Footprint for a single H36M inference (RTX 6000)